If 5 times the first number plus three times the second number eaguals 47, and 10 times the first number minus 4 times the second number equals 54, what are the numbers?
Answer by checkley75(3666)

5X+3Y=47
10X-4Y=54 NOW MULTIPLY THE FIRST EQUATION BY -2 & ADD:
-10X-6Y=-94
--------------------------
-10Y=-40
Y=-40/-10
Y=4 ANSWER
5X+3*4=47
5X+12=47
5X=47-12
5X=35
X=35/5
X=7 ANSWER.
PTOOF
10*7-4*4=54
70-16=54
54=54